OpenStudy (anonymous):

let T:P1->P2 be defined by T(p(x))=xP(x) Find the bases for Im(T) and Ker(T)

OpenStudy (anonymous):

@TuringTest help pls

OpenStudy (anonymous):

i'm not sure i did this! natural basis for P1 is {1,x^2} T(1)=x T(x)=x^2 Im(T)=<x,x^2> let a,b b element of R a=0 , b=0 T:P1->P2, T[P(x)]=x[P(x)] thus Im(T)=<X,X^2>

OpenStudy (anonymous):

@REMAINDER

OpenStudy (anonymous):

you are correct then dim(im(T)+dim(ker(T))=dim(P1) 2+dim(ker(T))=2 thus dim(ker(T))=0 then ker{0v}

OpenStudy (anonymous):

thanks a bunch neh, o tlalefile